My instuctions ask me to express each percent as a fraction or mixed number in simplest form, and as a decimal.
Please show me step by step how to solve 8 and 5 over 6% and also help me solve twenty-five and one over 3%

Answer by checkley77(12569)

I'M NOT EXACTLY SURE WHAT YOU WANT BUT i'LL TRY.
IF YOU MEAN 8/.06=133.333
IF YOU MEAN 5/.06=83.333
25/.03=833.333
1/.03=33.333:
OR DO YOU MEAN THE 'AND' TO BE A PLUS SIGN?
(8+5).06=13/.06=216.667
(25+1).03=26/.03=866.667